OpenStudy (juliebeans):

Expand the squared binomial. (x – 4)² A. x² − 16 B. x²− 8x + 16 C. x² + 16 D. x² + 8x − 16

OpenStudy (juliebeans):

@DivineSolar

OpenStudy (kendricklamar2014):

\[(x-4)^2 = (x+-4)(x+-4)\]

OpenStudy (juliebeans):

So..Is that what you simplfed to?

OpenStudy (kendricklamar2014):

yes, but you can simplify it even more

OpenStudy (kendricklamar2014):

\[(x+-4)(x+-4) = (x)(x) + (x)(-4)+(-4)(x)+(-4)(-4)\]

OpenStudy (juliebeans):

Can We Combine the x's on each side?

OpenStudy (kendricklamar2014):

\[= x^2 - 4x-4x+16\]

OpenStudy (kendricklamar2014):

What is:\[-4x - 4x = ?\]

OpenStudy (juliebeans):

Then combine 4x?

OpenStudy (kendricklamar2014):

Yes ^

OpenStudy (juliebeans):

Sorry That Was A lagg

OpenStudy (kendricklamar2014):

What is -4x - 4x

OpenStudy (juliebeans):

-8x?

OpenStudy (kendricklamar2014):

Correct

OpenStudy (juliebeans):

So Its B??

OpenStudy (kendricklamar2014):

So with that you get the answer:\[x^2 - 8x+16\]

OpenStudy (juliebeans):

Thanks!!!!

OpenStudy (kendricklamar2014):

No Problem

OpenStudy (juliebeans):

Can you help me with one more?

OpenStudy (kendricklamar2014):

Sure

OpenStudy (juliebeans):

@KendrickLamar2014 YASSSS NOW ITS WORKINGGG :D

OpenStudy (juliebeans):

11) Use FOIL to multiply the two binomials. (b – 3)(b + 9) A. b ² + 6b − 27 B. b ² − 27 C. b ² − 6b − 27 D. b ² + 12b – 27

OpenStudy (kendricklamar2014):

What do you think the answer is @JulieBeans

OpenStudy (juliebeans):

If I did my cominbing right i think its b again :P

OpenStudy (juliebeans):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

Nope

OpenStudy (kendricklamar2014):